y=(-x)/(2)-4
-2x-y=-5

Answer by stanbon(75887)

y=(-x)/(2)-4
-2x-y=-5
-----
Multiply thru the 1st equation by 2:
2y = -x -8
-2x -y = -5
-------------------
Rearrange:
x + 2y = -8
2x + y = 5
-----
Multiply thru the 1st equation by 2:
2x + 4y = -16
2x + y = 5
------
Subtract and solve for "y":
3y = -21
y = -7
--------
Solve for "x":
x = -2y-8
x = 14-8
x = 6
-----------
Solution: (6,-7)
=====================
cheers,
Stan H.